OpenStudy (anonymous):

Given that 2^(x) = 3^(y) =12^(z) , show that x = (z(2y+x))/y

OpenStudy (anonymous):

take log thru out the eqn gives x log2= y log3=zlog12 and log12=log 3 + log4 = log3 + 2log2 , log3= (x log2)/y thus x log2 = z log12 = z (log 3 + 2log2) = z( [(xlog2)/y ] +2 log2) thus x= z( [x/y]+2) =[z (x+2y) ]y

OpenStudy (anonymous):

thanks